A carpenter is making doors that are 2058 millimeters tall. If the doors are too long they must be trimmed, and if they are too short they cannot be used. A sample of 34 doors is taken, and it is found that they have a mean of 2048 millimeters. Assume a population variance of 441. Is there evidence at the 0.1level that the doors are too short and unusable?

Step 2 of 6:

Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

We are given

Xbar = 2048

σ^2 = 441

σ = sqrt(441) = 21

µ = 2058

n = 34

Test statistic formula is given as below:

Z = (Xbar - µ)/[σ/sqrt(n)]

Z = (2048 - 2058)/[21/sqrt(34)]

Z = -10/ 3.60147

Z = -2.77664

Test statistic = -2.78